Sol Summary – November 28th

Crew 183+ Sol Summary Report 28November2017

Sol 1

Title: Departure of Crew 183

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

This was the last day for Crew 183. They packed the last of their belongings and headed off, back to Earth. After eating breakfast, Shannon and I started assembling the new rovers. It took over an hour to get just one close to being ready to go, as the rovers didn’t come with assembly instructions. Unfortunately, we don’t have a socket wrench that is compatible with the large bolts used with the rovers. Hopefully we’ll be able to find a socket large enough to work.

At around 2:30 PM, we headed off to Cactus Rd, to rediscover a road that was thought to be lost and forgotten to time. Our goal was to see if this area could be accessed and explored by future EVAs. It was obvious it hadn’t been used by ATVs or rovers in years, as the most prominent tracks were from cows. The cow tracks made the road extremely bumpy, slowing our progress significantly.

Once we reached the end, we found a wonderful place to access the Summerville Formation, an easy 100 meter walk. There were these red and yellow pieces of chert-like rocks scattered everywhere that could be petrified dinosaur poop.

Once we finished here, we made our way to the end of Yellow Zebra road and found an amazing place for future EVAs. This area could easily take several hours to explore adequately. As it got closer to dark, we made our way back to the Hab, and I resumed my soil analysis in the Lab.

Look Ahead Plan:

Tomorrow, we plan to gather food supplies for the next incoming crew. I also intend to start exploring the region south of Kissing Camel range, to see if I find anything interesting. There is also a road in the area we conducted our EVA today that we still need to explore, called Maxwell Montes Rd. We may either do this tomorrow or sometime before Saturday, when Crew 184 arrives.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: End of Cactus Rd & Yellow Zebra Rd

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 28th

Crew 183 + Operations Report 28NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC  92% (Before generator is run at night)

Diesel –  50 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –  not known gallons.

Water (trailer) –  0 gallons.

Water (static) –   300 gallons

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 3.5

Notes on ATVs: nothing to report

Deimos rover used:  yes

Hours: not recorded (but crews need to record this information)

Beginning charge: 100

Ending charge: 95

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: We began putting the new rovers together
today.  I decided that we would construct the cage, but not modify the
seat at this time.  Instead, we will just remove the back seat rest,
which was what the Mars 160 crew wanted to do originally but I
insisted on the modification.  This will still be done on these but at
a later date because it is much more secure.  Maybe when Yusuke is
here with his crew, since he modified them last time and will be able
to help me to do it.  In the meantime, crews will still be able to use
them.  They will require about a month of short trips before they are
able to go on long ones like Deimos.

Summary of internet:  not known

Summary of suits and radios:  Ashton has recommended that all the old
radios that are sketchy be taken out of service.  I agree.  Four new
radios and headsets will be added to the Hab before the next crew
arrives.  Old sketchy ones will be removed.

Summary of Hab operations: Two new CO alarms were delivered to the
Hab.  We will mount one each on the lower and upper deck before the
new crew arrives.

Summary of GreenHab operations:  Two new planters were added to the
GreenHab.  They are for an herb garden and a greens garden.  I am
prepping the soil.  I turned the heater down to 55 degrees F.  Today I
left the door open all day.  Used 6 gallons of water for plants and

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support:  We have bolts
that have 10.8 and 8.8 on them.  What does that mean in turns of
socket size, if anything?  We don’t have a socket set that fits the
bolts on the rovers, so are trying to figure out if they are metric or
English system.  Any information appreciated.

Copyright © The Mars Society. All rights reserved. | Main Site